# A Gaussian Process View on Observation Noise and Initialization in Wide Neural Networks

### v2

@@ 4.2 NTK-GP Posterior Mean with Aleatoric Noise Interpretation
It is important to highlight that the trained parameters $\theta_{\infty}$ generally depend on the specific initialization $\theta_{0}$, and as such, will differ for every random initialization. This variability is relevant to the training process but is not central to our analysis and use case. Unlike deep ensembles (he2020bayesian) (particularly when used in the context of Thompson sampling (thompson-sampling)), where the distribution of trained networks across initialization is explicitly used, our focus lies on the behaviour of a single trained network under a given initialization. We are able to alleviate the effect of this randomness thanks to our initialization strategy described in the next section.
@@ 5 Neural Network Initialization as NTK-GPs with Arbitrary Prior Mean
While standard initialization schemes for neural networks typically involve randomly setting weights and biases to break symmetry, they generally lack a principled way to align the network’s prior outputs with specific inductive biases or desired properties. In contrast, Gaussian processes offer a well-defined framework for specifying prior distributions over functions through the prior mean and covariance. Arbitrary initialization is especially valuable in reinforcement learning, where optimistic priors aid exploration (strehl2006pac), and in domains with strong prior knowledge, where encoding a meaningful prior mean can improve performance.
One way to use the results presented in the previous section to obtain the posterior of the NTK-GP with zero prior mean is to train several ensembles and average the outputs using the law of large numbers, akin to the approach by he2020bayesian. However, if we are only interested in computing the posterior mean and not the covariance, there is a better way, which only requires the use of one network. Section 5.1 explores how neural networks, under the NTK-GP framework, can be initialized to reflect arbitrary prior means, enhancing their flexibility to align with specific tasks. Appendix B shows why simply setting the last layer’s weights to zero does not yield the desired prior.
@@ 5.1 Shifting the labels or predictions
Inspired by standard techniques in GP literature (williams2006gaussian), we provide a formal construction for modifying the network to introduce arbitrary prior mean. This can be accomplished either by shifting the predictions of the neural network by its predictions at initialization, or equivalently, defining a new shifted network. The following theorem formalizes this, resolving the main research question we posed in Section 3.
@@
We prove this in Appendix H. Note that $\tilde{f}_{\theta_{0}}(\mathbf{x^{\prime}},\theta_{\infty})$ is non-random (i.e. independent of the initialization $\theta_{0}$). This is different from training the non-shifted network, where $f_{\theta_{0}}(\mathbf{x^{\prime}},\theta_{\infty})$ is random, but has a mean which corresponds to the posterior mean of an NTK-GP with prior mean 0. Note, that the non-randomness of $\tilde{f}_{\theta_{0}}(\mathbf{x^{\prime}},\theta_{\infty})$ is not immediately clear: while $\tilde{f}_{\theta_{0}}(x,\theta_{0})$ does not depend on $\theta_{0}$, $\tilde{f}_{\theta_{0}}(x,\theta)$ differs for different $\theta_{0}$ when $\theta\neq\theta_{0}$. The reason for the non-randomness after training is that the NTK doesn’t depend on the initialization in the infinite-width limit.
@@ 6.1 Experimental Setup
To ensure alignment with the NTK framework, we train wide fully connected multi-layer perceptrons (MLPs) under the NTK parametrization using full-batch gradient descent (additional results under SGD with different batch sizes are included in Appendix C). The experiments are performed in both synthetic and standard regression datasets. The synthetic problem is a multi-dimensional synthetic regression task where the target function is defined as $\mathbf{y}=\sin(\mathbf{x})+\cos(2\mathbf{x})+\boldsymbol{\epsilon}$, with $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I\sigma^{2})$ representing the observation noise. Inputs $\mathbf{x}$ are uniformly sampled from the range [6,6][-6,6]. We generate a dataset of a few hundred training points and use a two-layer neural network. For the standard regression tasks, we use the UCI datasets Airline, Taxi, and Year, which are complex and high-dimensional. We follow bergna2025post for architectural choices and hyperparameter settings. All experiments were performed on a NVIDIA L40 GPU.
@@
Finally, in Section 6.4, we explore how incorporating a learned prior mean can improve data efficiency in transfer settings. We compare training from scratch to initializing with the posterior from a related task, and observe significantly improved performance, especially in low-data regimes. We will release the complete code on acceptance, including a self-contained Jupyter notebook for reproducibility.
@@ 6.2 Empirical Convergence of the Parameters
Figure 1: Frobenius norm differences between the trained neural network’s parameters and the kernel ridge regression solution plotted against network width for different datasets. Left: Different input dimensions of the synthetic dataset. Middle: Airline and Taxi UCI datasets. Right: Year UCI dataset. Shaded regions represent the standard deviation divided by the square root of the number of seeds.
@@ 6.3 Empirical Convergence of the Trained Network to a Linear Model
Figure 2: Supremum of the $\ell_{2}$ norm between the trained neural network’s outputs $f(x,\theta_{\infty})$ and the linearized model’s predictions $f^{\text{lin}}(x,\theta_{\infty}^{\text{lin}})$ across the validation set, plotted against network width. Left: Different input dimensions of the synthetic dataset. Middle: Airline and Taxi UCI datasets. Right: Year UCI dataset.
@@ 6.4 Pre-Trained Neural Network as Prior Mean
In this experiment, we construct two $N$-dimensional regression tasks that share low-frequency structure but differ in their high-frequency components. In both cases, observations follow $\mathbf{y}=f(\mathbf{x})+\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I\sigma^{2})$. Task 1 uses the target function $f_{1}(\mathbf{x})=\sin(\mathbf{x})+0.5\sin(5\mathbf{x})+0.2\sin(20\mathbf{x})$, while Task 2 is defined by $f_{2}(\mathbf{x})=\sin(\mathbf{x})+0.3\cos(7\mathbf{x})-0.2\sin(15\mathbf{x})$. Coefficients were chosen arbitrarily and were not tuned.
@@ 7 Related Work
lee2019wide demonstrates that as network width approaches infinity, training dynamics simplify and can be approximated by a linearized model using a first-order Taylor expansion. lee2019wide also study the links between the output of trained neural networks and GPs. Crucially, we extend this work by proving that this linearization still holds in the presence of observation noise.
@@
jacot2018neural established the equivalence between wide neural networks and kernel methods via kernel ridge regression, and popularized studying neural networks in function space rather than parameter space. We build on both ideas: Theorem 5.1 adopts the function-space view of GPs. Subsequent work has further explored NTK–GP connections in noisy settings. For instance, rudner2023functionspaceregularizationneuralnetworks proposed function-space regularization to encode desired properties into predictions, while chen2022neural related NTK norms to RKHS regularization.
@@
In overparameterized settings, allenzhu2019convergence show that SGD can reach global minima in polynomial time, while zou2020gradient prove convergence for ReLU networks by keeping weights close to their initialization. Although we do not directly rely on these results, we also guarantee high-probability convergence to the global optimum.
@@ 8 Conclusion and Future Work
We studied the relationship between regularized training in wide neural networks and Gaussian Processes under the NTK framework. We are the first work formally showing that the proposed weight-space regularization in neural networks is equivalent to adding aleatoric noise to the NTK-GP posterior mean in the infinite-width limit. Additionally, we introduced a shifted network approach that enables arbitrary prior functions and ensures deterministic convergence to the NTK-GP posterior mean without requiring ensembles or kernel inversion. Empirical results validate our theoretical findings, demonstrating convergence to the linearized network. The methods proposed in this paper establish a foundation for practical Gaussian Process inference using the NTK and standard neural network training. Future work could explore the benefits of leveraging the NTK to efficiently train GPs with interpretable and widely used kernels, such as RBF or Matérn. This would require looking into algorithms to compute the posterior covariance (calvo2024epistemic; zanger2025contextualsimilaritydistillationensemble).
@@ Appendix A Local Boundedness and Lipschitzness of the Jacobian
To analyze the training dynamics of wide neural networks, it is essential to ensure that the Jacobian of the network remains (locally) bounded and Lipschitz-continuous during training, with the constants scaling “well” with layer width $n$. This is shown by the following Lemma, which was proved in lee2019wide for standard parametrization.
@@
$\forall\theta,\tilde{\theta}\in B(\theta_{0},C):\lVert J(x,\theta)-J(x,\tilde{\theta})\rVert_{2}\leq\frac{1}{\sqrt{n}}K^{\prime}\lVert\theta-\tilde{\theta}\rVert_{2}.$ (21)
@@
$\forall\theta,\tilde{\theta}\in B(\theta_{0},C):\lVert J(\theta)-J(\tilde{\theta})\rVert_{F}\leq\frac{1}{\sqrt{n}}K\lVert\theta-\tilde{\theta}\rVert_{2}.$ (23)
@@
$\forall\theta\in B(\theta_{0},C):\left\lVert\nabla^{2}_{\theta}f(x,\theta)\right\rVert_{2}\leq\frac{1}{\sqrt{n}}K^{\prime}.$ (24)
Here, the Frobenius norm is used to aggregate over different training points, i.e. $\lVert J(\theta)\rVert_{F}^{2}=\sum_{i=1}^{N}\lVert J(\mathbf{x}_{i},\theta)\rVert_{2}^{2}$.
lee2019wide proved this Lemma for standard parametrization. In their version for NTK parametrization (Lemma 2), there is unfortunately a typo stating that the Lipschitz constant of the Jacobian is $O(1)$ instead of $O(\frac{1}{\sqrt{n}})$, which is why we quickly go over how to get this Lemma. Note that we do not train the first layer and the biases for simplicity. Let $f^{\operatorname*{ntk}}$ be the network in NTK parametrization, and let $\theta_{0}$ be randomly initialized according to the NTK parametrization. Further, let $\theta,\tilde{\theta}\in B(\theta_{0},C)$. Then, $f^{\operatorname*{std}}(\theta):=f^{\operatorname*{ntk}}(\sqrt{n}\theta)$ is a network in standard parametrization, and $\frac{1}{\sqrt{n}}\theta_{0}$ is randomly initialized in standard parametrization. Further, $\frac{1}{\sqrt{n}}\theta,\frac{1}{\sqrt{n}}\tilde{\theta}\in B(\frac{1}{\sqrt{n}}\theta_{0},\frac{1}{\sqrt{n}}C)$. Note, that $J^{\operatorname*{ntk}}(\theta)=\frac{1}{\sqrt{n}}J^{\operatorname*{std}}(\frac{1}{\sqrt{n}}\theta)$. Now, by applying Lemma 1 from lee2019wide to the network in standard parametrization and the parameters $\frac{1}{\sqrt{n}}\theta_{0},\frac{1}{\sqrt{n}}\theta,\frac{1}{\sqrt{n}}\tilde{\theta}$, we get with high probability over random initialization:
$\lVert J^{\operatorname*{ntk}}(x,\theta)\rVert_{2}=\frac{1}{\sqrt{n}}\lVert J^{\operatorname*{std}}(x,\frac{1}{\sqrt{n}}\theta)\rVert_{2}\leq\frac{1}{\sqrt{n}}\sqrt{n}K^{\prime}=K^{\prime},$ (25)
and
$\lVert J^{\operatorname*{ntk}}(x,\theta)-J^{\operatorname*{ntk}}(x,\theta)\rVert_{2}$ $=\frac{1}{\sqrt{n}}\lVert J^{\operatorname*{std}}(x,\frac{1}{\sqrt{n}}\theta)-J^{\operatorname*{std}}(x,\frac{1}{\sqrt{n}}\tilde{\theta})\rVert_{2}$ (26)
$\leq\frac{1}{\sqrt{n}}\sqrt{n}K^{\prime}\lVert\frac{1}{\sqrt{n}}\theta-\frac{1}{\sqrt{n}}\tilde{\theta}\rVert_{2}=\frac{1}{\sqrt{n}}K^{\prime}\lVert\theta-\tilde{\theta}\rVert_{2}.$ (27)
This result establishes that the Jacobian and its changes remain controlled during training, with the Lipschitz constant scaling appropriately with the network width.
Appendix B Initializing the last layer weights as 0 does not lead to a zero prior mean
@@
This actually converges to the Neural Network Gaussian Process (NNGP) kernel (lee2018dnnsgps), which does not align with the desired result.
However, our observation highlights that training the network in this configuration corresponds to performing inference in a Gaussian Process (GP) with a zero prior mean and the NNGP kernel as the covariance function. Specifically, in the case of the linearized network (or a network with a sufficiently large layer width), the partial derivatives with respect to the hidden layer parameters are zero. Thus, effectively only the last layer is trained. As shown in lee2019wide, this scenario is equivalent to performing inference with the NNGP, where the covariance of the trained network matches the posterior covariance of the NNGP.
@@ Appendix D Regularized gradient flow and gradient descent for the linearized network
Consider the linearized network $f^{\operatorname*{lin}}(x,\theta)=f(x,\theta_{0})+J(x,\theta_{0})(\theta-\theta_{0})$. For notational convenience, we drop the dependence on $\theta_{0}$ throughout the Appendix. In the following, we will consider training the parameters using the regularized training loss

### v3

@@ 4.2 NTK-GP Posterior Mean with Aleatoric Noise Interpretation
It is important to highlight that the trained parameters $\theta_{\infty}$ generally depend on the specific initialization $\theta_{0}$, and as such, will differ for every random initialization. This variability is relevant to the training process but is not central to our analysis and use case. Unlike deep ensembles (He et al., 2020) (particularly when used in the context of Thompson sampling (Thompson, 1933)), where the distribution of trained networks across initialization is explicitly used, our focus lies on the behaviour of a single trained network under a given initialization. We are able to alleviate the effect of this randomness thanks to our initialization strategy described in the next section.
@@ 5 Neural Network Initialization as NTK-GPs with Arbitrary Prior Mean
This section addresses limitation (ii), arbitrary prior means, by introducing a shifted network construction that enables deterministic NTK-GP posterior means with user-specified priors. While standard initialization schemes for neural networks typically involve randomly setting weights and biases to break symmetry, they generally lack a principled way to align the network’s prior outputs with specific inductive biases or desired properties. In contrast, Gaussian processes offer a well-defined framework for specifying prior distributions over functions through the prior mean and covariance. Arbitrary initialization is especially valuable in reinforcement learning, where optimistic priors aid exploration (Strehl et al., 2006), and in domains with strong prior knowledge, where encoding a meaningful prior mean can improve performance.
One way to use the results presented in the previous section to obtain the posterior of the NTK-GP with zero prior mean is to train several ensembles and average the outputs using the law of large numbers, akin to the approach by He et al. (2020). However, if we are only interested in computing the posterior mean and not the covariance, there is a better way, which only requires the use of one network. Section 5.1 explores how neural networks, under the NTK-GP framework, can be initialized to reflect arbitrary prior means, enhancing their flexibility to align with specific tasks. Appendix B shows why simply setting the last layer’s weights to zero does not yield the desired prior.
@@ 5.1 Shifting the labels or predictions
Inspired by standard techniques in GP literature (Williams and Rasmussen, 2006), we provide a formal construction for modifying the network to introduce arbitrary prior mean. This can be accomplished either by shifting the predictions of the neural network by its predictions at initialization, or equivalently, defining a new shifted network. The intuition behind this is that the shifting does not change the Jacobian of the network with respect to the parameters, but gives a different initialization in Equation (11). The following theorem formalizes this, resolving the main research question we posed in Section 3.
@@
We prove this in Appendix H. Note that $\tilde{f}_{\theta_{0}}(\mathbf{x^{\prime}},\theta_{\infty})$ is non-random (i.e., independent of the initialization $\theta_{0}$). This is different from training the non-shifted network, where $f_{\theta_{0}}(\mathbf{x^{\prime}},\theta_{\infty})$ is random, but has a mean which corresponds to the posterior mean of an NTK-GP with prior mean 0. Note, that the non-randomness of $\tilde{f}_{\theta_{0}}(\mathbf{x^{\prime}},\theta_{\infty})$ is not immediately clear: while $\tilde{f}_{\theta_{0}}(x,\theta_{0})$ does not depend on $\theta_{0}$, $\tilde{f}_{\theta_{0}}(x,\theta)$ differs for different $\theta_{0}$ when $\theta\neq\theta_{0}$. The reason for the non-randomness after training is that the NTK doesn’t depend on the initialization in the infinite-width limit.
@@ 6.1 Experimental Setup
Figure 1: Frobenius norm differences between the trained neural network’s parameters and the kernel ridge regression solution plotted against network width for different datasets. Left: Different input dimensions of the synthetic dataset. Middle: Airline and Taxi datasets. Right: UCI Year dataset. Shaded regions represent the standard deviation divided by the square root of the number of seeds.
Figure 2: Supremum of the $\ell_{2}$ norm between the trained neural network’s outputs $f(x,\theta_{\infty})$ and the linearized model’s predictions $f^{\text{lin}}(x,\theta_{\infty}^{\text{lin}})$ across the validation set, plotted against network width. Left: Different input dimensions of the synthetic dataset. Middle: Airline and Taxi datasets. Right: UCI Year dataset.
To ensure alignment with the NTK framework, we train wide fully connected multi-layer perceptrons (MLPs) under the NTK parametrization using full-batch gradient descent (additional results under SGD with different batch sizes are included in Appendix C). The experiments are performed in both synthetic and standard regression datasets. The synthetic problem is a multi-dimensional synthetic regression task where the target function is defined as $\mathbf{y}=\sin(\mathbf{x})+\cos(2\mathbf{x})+\boldsymbol{\epsilon}$, with $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I\sigma^{2})$ representing the observation noise. Inputs $\mathbf{x}$ are uniformly sampled from the range [6,6][-6,6]. We generate a dataset of a few hundred training points and use a two-layer neural network. For the standard regression tasks, we use the datasets Airline (Hensman et al., 2013), Taxi (Peng et al., 2017; Salimbeni and Deisenroth, 2017), and UCI Year (Hernández-Lobato and Adams (2015)), which are complex and high-dimensional. We follow Bergna et al. (2025) for architectural choices and hyperparameter settings. All experiments were performed on a single NVIDIA L40 GPU.
@@
Finally, in Section 6.4, we explore how incorporating a learned prior mean can improve data efficiency in transfer settings. We compare training from scratch to initializing with the posterior from a related task, and observe significantly improved performance, especially in low-data regimes.
@@ 6.4 Pre-Trained Neural Network as Prior Mean
In this experiment, we construct two $N$-dimensional regression tasks that share low-frequency structure but differ in their high-frequency components. In both cases, observations follow $\mathbf{y}=f(\mathbf{x})+\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\sigma^{2}I)$. Task 1 uses the target function $f_{1}(\mathbf{x})=\sin(\mathbf{x})+0.5\sin(5\mathbf{x})+0.2\sin(20\mathbf{x})$, while Task 2 is defined by $f_{2}(\mathbf{x})=\sin(\mathbf{x})+0.3\cos(7\mathbf{x})-0.2\sin(15\mathbf{x})$. Coefficients were chosen arbitrarily and were not tuned.
@@ 7 Related Work
Lee et al. (2019) demonstrate that as network width approaches infinity, training dynamics simplify and can be approximated by a linearized model using a first-order Taylor expansion. Lee et al. (2019) also study the links between the output of trained neural networks and GPs. Crucially, we extend this work by proving that this linearization still holds in the presence of observation noise.
@@
Jacot et al. (2018) established the equivalence between wide neural networks and kernel methods via kernel ridge regression, and popularized studying neural networks in function space rather than parameter space. We build on both ideas: Theorem 5.1 adopts the function-space view of GPs. Subsequent work has further explored NTK–GP connections in noisy settings. For instance, Rudner et al. (2023) proposed function-space regularization to encode desired properties into predictions, while Chen et al. (2022) related NTK norms to RKHS regularization.
@@
In overparameterized settings, Allen-Zhu et al. (2019) show that SGD can reach global minima in polynomial time, while Zou et al. (2020) prove convergence for ReLU networks by keeping weights close to their initialization. Although we do not directly rely on these results, we also guarantee high-probability convergence to the global optimum.
@@ 8 Conclusion and Future Work
We studied the relationship between regularized training in wide neural networks and Gaussian Processes under the NTK framework. We are the first work formally showing that the proposed weight-space regularization in neural networks is equivalent to adding aleatoric noise to the NTK-GP posterior mean in the infinite-width limit. Additionally, we introduced a shifted network approach that enables arbitrary prior functions and ensures deterministic convergence to the NTK-GP posterior mean without requiring ensembles or kernel inversion. Empirical results validate our theoretical findings, demonstrating convergence to the linearized network. The methods proposed in this paper establish a foundation for practical Gaussian Process inference using the NTK and standard neural network training. Future work could explore the benefits of leveraging the NTK to efficiently train GPs with interpretable and widely used kernels, such as RBF or Matérn. This would require looking into algorithms to compute the posterior covariance (Calvo-Ordoñez et al., 2024; Zanger et al., 2025).
Acknowledgements
SCO’s research is supported by the Oxford-Man Institute through the EPSRC Centre for Doctoral Training in Mathematics of Random Systems: Analysis, Modelling and Simulation (EPSRC Grant EP/S023925/1). JP acknowledges financial support from the Oxford-Man Institute. SCO, JP, and AC acknowledge the support of the Oxford-Man Institute for providing computational resources. JMHL acknowledges support from EPSRC funding under grant EP/Y028805/1. JMHL also acknowledges support from a Turing AI Fellowship under grant EP/V023756/1.
@@ Appendix A Local Boundedness and Lipschitzness of the Jacobian
To analyze the training dynamics of wide neural networks, it is essential to ensure that the Jacobian of the network remains (locally) bounded and Lipschitz-continuous during training, with the constants scaling “well” with layer width $n$. This is shown by the following Lemma, which was proved in Lee et al. (2019) for standard parametrization, and more formally in Liu et al. (2020) for NTK parameterization.
@@
$\forall\theta,\tilde{\theta}\in B(\theta_{0},C):\lVert J(x,\theta)-J(x,\tilde{\theta})\rVert_{2}\leq\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}\lVert\theta-\tilde{\theta}\rVert_{2}.$ (21)
@@
$\forall\theta,\tilde{\theta}\in B(\theta_{0},C):\lVert J(\theta)-J(\tilde{\theta})\rVert_{F}\leq\frac{(\log n)^{c}}{\sqrt{n}}K\lVert\theta-\tilde{\theta}\rVert_{2}.$ (23)
@@
$\forall\theta\in B(\theta_{0},C):\left\lVert\nabla^{2}_{\theta}f(x,\theta)\right\rVert_{2}\leq\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}.$ (24)
The Frobenius norm is used to aggregate over different training points, i.e. $\lVert J(\theta)\rVert_{F}^{2}=\sum_{i=1}^{N}\lVert J(\mathbf{x}_{i},\theta)\rVert_{2}^{2}$ for $J(\theta)=J(\mathbf{x},\theta)\in\mathbb{R}^{N\times p}$. The logarithmic term $(\log n)^{c}$ is sometimes omitted in the literature. Its power $c>0$ is a constant that only depends on the network architecture.
Appendix B Initializing the last layer weights as 0 does not lead to a zero prior mean NTK-GP
@@
This actually converges to the Neural Network Gaussian Process (NNGP) kernel (Lee et al., 2018), which does not align with the desired result.
However, our observation highlights that training the network in this configuration corresponds to performing inference in a Gaussian Process (GP) with a zero prior mean and the NNGP kernel as the covariance function. Specifically, in the case of the linearized network (or a network with a sufficiently large layer width), the partial derivatives with respect to the hidden layer parameters are zero. Thus, effectively only the last layer is trained. As shown in Lee et al. (2019), this scenario is equivalent to performing inference with the NNGP, where the covariance of the trained network matches the posterior covariance of the NNGP.
@@ Appendix C Additional Experiments
Method Test RMSE $\downarrow$
Exact Matern GP OOM
Inducing Points Matern GP (200) 0.075
NTK-GP via SGD (ours) 0.099
Ensemble of 5 Wide MLPs 0.104
Wide MLP (no regularization) 0.788
Shallow MLP 0.201
Table 1: Comparison on the Kin8nm regression dataset. We compare our method (NTK-GP posterior mean via SGD with zero prior mean) against exact and approximate Matern GPs, ensembles of wide MLPs (He et al., 2020), an unregularized wide MLP (Lee et al., 2019), and a shallow MLP. Exact GP inference resulted in out-of-memory (OOM) issues. Our method outperforms the ensemble while requiring only a single training run and provides a scalable approximation to the GP posterior mean.
@@ Appendix D Regularized gradient flow and gradient descent for the linearized network
Consider the linearized network $f^{\operatorname*{lin}}_{\theta_{0}}(x,\theta)=f(x,\theta_{0})+J(x,\theta_{0})(\theta-\theta_{0})$. For notational convenience, we drop the dependence on $\theta_{0}$ throughout the Appendix. In the following, we will consider training the parameters using the regularized training loss